OpenStudy (impracticaljoker):

(02.01 LC) Solve: The quantity 2 x minus 20 divided by 3 = 2x (5 points) −5 −3 2 13 @quickstudent I am in need of assistance

jhonyy9 (jhonyy9):

(2x-20)/3 = 2x do you think hence ?

OpenStudy (deathlynine):

-5

OpenStudy (deathlynine):

http://openstudy.com/updates/55b811c7e4b0c7a872f8706e This one explains it well

jhonyy9 (jhonyy9):

Deathlynine how you get this result ? please just help to asker the way how can solving same exercise not directly answer never please ! ty.

OpenStudy (deathlynine):

@jhonyy9 show her the step then.

jhonyy9 (jhonyy9):

sorry but you begined started the help with this answer of 5 - how you get this result ?

jhonyy9 (jhonyy9):

joker i like please you collaborate - ok. ?

jhonyy9 (jhonyy9):

bc. on the left side there is a fraction what you need to do first of all ?

jhonyy9 (jhonyy9):

do you know what mean common denominator ?

jhonyy9 (jhonyy9):

or we can rewriting these all in this form (2x-20) 2x ------- = ---- 3 1 how you like it now ?

jhonyy9 (jhonyy9):

ok. so than you cross multiply what will get ?

jhonyy9 (jhonyy9):

joker please collaborate - i not can give you directly the right answer

OpenStudy (impracticaljoker):

I had to help my mother do something. I was even on the computer.

jhonyy9 (jhonyy9):

so than you cross multiply will get 1(2x-20) = 3*2x what mean 2x-20=3*2x do you can continue from here ?

OpenStudy (impracticaljoker):

Yes. I'll just follow fromhere

jhonyy9 (jhonyy9):

ok. please collaborate from here than - ty.

OpenStudy (impracticaljoker):

Ok so where do you want me to start?

jhonyy9 (jhonyy9):

from here how you get the value of x ? 2x-20=3*2x

OpenStudy (impracticaljoker):

Wouldn't I add 20 to both sides first?

OpenStudy (quickstudent):

(2x - 20)/3 = 2x Is this what the equation really looks like?

jhonyy9 (jhonyy9):

quickstudent please give me possible help - thank you

jhonyy9 (jhonyy9):

yes joker this is a right step

OpenStudy (impracticaljoker):